# EvoEmo: Towards Evolved Emotional Policies for Adversarial LLM Agents in Multi-Turn Price Negotiation

### v3

@@ Abstract.
Recent research on Chain-of-Thought (CoT) reasoning in Large Language Models (LLMs) has demonstrated their capability for complex, multi-turn negotiations, a task that inherently involves understanding and leveraging human emotional cues. As autonomous LLM agents are increasingly deployed to perform such tasks, a new paradigm of LLM-vs-LLM interaction is emerging as a critical research domain. However, while these agents are built on models trained to process human emotion, existing negotiation strategies for LLM agents largely overlook the functional role of emotions as a strategic action. This renders them passive and vulnerable to manipulation and strategic exploitation by other, more sophisticated LLM counterparts. To address this gap, we present EvoEmo, an evolutionary reinforcement learning framework that optimizes dynamic emotional expression in negotiations. EvoEmo models emotional state transitions as a Markov Decision Process and employs population-based genetic optimization to evolve high-reward emotion policies across diverse negotiation scenarios. We further propose an evaluation framework with two baselines—vanilla strategies and fixed-emotion strategies—for benchmarking emotion-aware negotiation. Extensive experiments and ablation studies show that EvoEmo consistently outperforms both baselines, achieving higher success rates, higher efficiency, and increased buyer savings. This findings highlight the importance of adaptive emotional expression in enabling more effective LLM agents for multi-turn negotiation.
@@ 1. Introduction
Extensive behavioral research has established that human decision-making deviates from classical economic rationality, being shaped by psychological biases and emotional states (Hilbert, 2012; Baumeister et al., 2012; Riaz et al., 2012). While modern Large Language Models (LLMs) have made progress in replicating personality-driven behaviors (Wei et al., 2025), the role of emotion as a strategic, dynamic force remains understudied (Liu and Long, 2025), especially compared to static, trait-based approaches (Huang and Hadfi, 2024). This gap is particularly critical in fine-grained negotiation scenarios such as price bargaining (Lin et al., 2023), where emotions directly influence on tactical choices (e.g., frustration prompting premature concessions, or excitement triggering overly aggressive bids), with immediate consequences for negotiation outcomes.
However, a paradigm shift is underway. Compared to the traditional focus on LLM agent-human user interaction, autonomous LLM agents are increasingly being deployed to perform complex tasks like negotiation amongst themselves (Lin et al., 2024). The natural consequence of this proliferation is a new ecosystem defined by agent-to-agent interactions (Abbasiantaeb et al., 2024; Lin et al., 2024). In domains such as e-commerce, supply chain management, and decentralized autonomous organizations (DAOs) (Hu et al., 2025), the scale of transactions necessitates agents that can negotiate with each other directly, with minimal human intervention. This emerging “agentic ecology” demands a focus on strategic robustness, where an agent’s ability to avoid exploitation and achieve favorable outcomes is paramount. The strategic value of emotion in this LLM-vs-LLM context stems not from anthropological fidelity but from the foundational architecture of the underlying language model. An LLM’s next-token prediction objective, trained on a corpus of human communication, has implicitly encoded emotional cues as high-dimensional features that strongly condition model output. Thus, emotional labels serve as a deterministic steering mechanism within the model’s probabilistic generation space where emotional cues (e.g., “frustration,” “satisfaction”) are strongly correlated with specific stances and potential actions. This makes an emotional signal a available control mechanism, which can directly shape an opponent LLM’s response generation.
@@
Despite this, current LLM agents are fundamentally limited in their use of emotion. Techniques like Direct Preference Optimization (DPO) (Gao et al., 2025) and Reinforcement Learning from Human Feedback (RLHF) (Kasbouya and Sael, 2025) have made them proficient in recognizing and reacting to human emotional cues, but these capabilities remain reactive. They fail to leverage emotion proactively as a tool to influence the negotiation trajectory and assert strategic dominance. This leaves them vulnerable in adversarial LLM-vs-LLM interactions, where they exhibit three critical deficiencies:
@@ 2.1. Emotions in Negotiation Dynamics
While personality traits influence general negotiation tendencies (e.g., agreeableness correlating with cooperative strategies) (Huang and Hadfi, 2024), emotions dominate real-time bargaining dynamics through three key mechanisms. First, temporal alignment ensures that negotiations—unfolding over rapid conversational turns lasting seconds to minutes—operate on a timescale closely aligned with emotional fluctuations, such as facial expressions and transient mood shifts, thereby often overriding longer-term personality influences (Griessmair et al., 2015). Second, strategic flexibility allows negotiators to adapt emotional expressions contextually; for instance, displaying tactical anger to convey firmness or simulate scarcity, temporarily overriding personality defaults (cite). Third, interactive amplification describes the process of emotional contagion, where one party’s affective stance elicits corresponding emotions in the opponent, forming feedback loops that modulate bargaining power over multiple turns (Olekalns and Druckman, 2014). This fine-grained affective interplay underpins dynamic shifts in human negotiation behavior, making emotions a decisive factor in real-time decision-making.
@@ 2.2. LLM Agents in Negotiation Systems
Current LLM-based negotiation systems demonstrate significant capabilities in emotion recognition but fall short in strategic emotional adaptation. Frameworks like AgreeMate (Chatterjee et al., 2024) and ACE (Shea et al., 2024) show proficiency in inferring emotions through chain-of-thought reasoning, yet lack mechanisms for dynamic emotional strategy evolution. For instance, AgreeMate’s modular architecture optimizes fixed buyer/seller roles but cannot evolve emotional strategies amid negotiation. Similarly, ACE’s coaching system detects errors but cannot simulate the real-time emotional volatility that characterizes human bargaining. Consequently, current LLM agents remain largely reactive rather than proactive in their negotiation behavior.
Recent reinforcement learning advances have improved LLM negotiation capabilities. Methods such as Q-learning (Watkins and Dayan, 1992) and PPO (Schulman et al., 2017) have been adapted to optimize negotiation policies via reward maximization. GENTEEL-NEGOTIATOR (Priya et al., 2025) further advances polite negotiation dialogue systems through mixture-of-expert reinforcement learning with specialized rewards for strategy, politeness, and coherence. Multi-agent evolutionary algorithms like EvoAgent (Yuan et al., 2024) offer promising pathways for strategy enhancement through population-based optimization. However, these approaches face fundamental limitations in emotion-aware negotiation scenarios. Reinforcement learning methods struggle with the sparse-reward, high-volatility nature of negotiation environments, while evolutionary algorithms typically evaluate policies based on terminal outcomes, providing weak credit assignment for the nuanced interplay between consecutive emotional states in multi-turn negotiations. Crucially, these methods lack online learning capabilities essential for price negotiation across diverse products and scenarios. Real-world negotiation agents must adapt emotional strategies in real-time when facing different product types, price ranges, and debtor circumstances without requiring separate training for each variation. Our approach integrates evolutionary exploration with reinforcement learning to enable online emotion policy optimization, allowing agents to continuously adapt emotional strategies during execution across varied product negotiations.
@@ 4.1. Datasets
Likewise to Huang and Hadfi (2024), we selected a subset of negotiation cases from the CraigslistBargain dataset (He et al., 2018)—a widely-used benchmark for negotiation studies—for evaluation. This subset comprises 20 distinct multi-turn negotiation scenarios spanning diverse product categories, including electronics, furniture, vehicles, and housing. Each scenario specifies: (1) product details (name, category, and description), (2) agent-specific target prices for the seller, and (3) emotion annotations that support realistic bargaining dynamics. The selected dataset spans a wide price range (\$50–\$5,000) and includes varied item conditions (brand new or used), thereby enabling comprehensive evaluation of negotiation policies under heterogeneous market conditions.

### v4

@@ Abstract.
Recent research on Chain-of-Thought (CoT) reasoning in Large Language Models (LLMs) has demonstrated their capability for complex, multi-turn negotiations, a task that inherently involves understanding and leveraging human emotional cues. As autonomous LLM agents are increasingly deployed to perform such tasks, a new paradigm of LLM-vs-LLM interaction is emerging as a critical research domain. However, while these agents are built on models trained to process human emotion, existing negotiation strategies for LLM agents largely overlook the functional role of emotions as a strategic action. This renders them passive and vulnerable to manipulation and strategic exploitation by other, more sophisticated LLM counterparts. To address this gap, we present EvoEmo, an evolutionary reinforcement learning framework that optimizes dynamic emotional expression in negotiations. EvoEmo models emotional state transitions as a Markov Decision Process and employs population-based genetic optimization to evolve high-reward emotion policies across diverse negotiation scenarios. We further propose an evaluation framework with two baselines—vanilla strategies and fixed-emotion strategies—for benchmarking emotion-aware negotiation. Extensive experiments and ablation studies show that EvoEmo consistently outperforms both baselines, achieving higher success rates, higher efficiency, and increased buyer savings. This findings highlight the importance of adaptive emotional expression in enabling more effective LLM agents for multi-turn negotiation. The code is available at https://github.com/Yunbo-max/EvoEmo.
@@ 1. Introduction
Extensive behavioral research has established that human decision-making deviates from classical economic rationality, being shaped by psychological biases and emotional states (hilbert2012toward; baumeister2012emotional; riaz2012personality). While modern Large Language Models (LLMs) have made progress in replicating personalitAMASS y-driven behaviors (wei2025mecot), the role of emotion as a strategic, dynamic force remains understudied (liu2025eq), especially compared to static, trait-based approaches (huang2024personality). This gap is particularly critical in fine-grained negotiation scenarios such as price bargaining (lin2023toward), where emotions directly influence on tactical choices (e.g., frustration prompting premature concessions, or excitement triggering overly aggressive bids), with immediate consequences for negotiation outcomes.
However, a paradigm shift is underway. Compared to the traditional focus on LLM agent-human user interaction, autonomous LLM agents are increasingly being deployed to perform complex tasks like negotiation amongst themselves (lin2024large). The natural consequence of this proliferation is a new ecosystem defined by agent-to-agent interactions (abbasiantaeb2024let; lin2024large). In domains such as e-commerce, supply chain management, and decentralized autonomous organizations (DAOs) (hu2025trustless), the scale of transactions necessitates agents that can negotiate with each other directly, with minimal human intervention. This emerging “agentic ecology” demands a focus on strategic robustness, where an agent’s ability to avoid exploitation and achieve favorable outcomes is paramount. The strategic value of emotion in this LLM-vs-LLM context stems not from anthropological fidelity but from the foundational architecture of the underlying language model. An LLM’s next-token prediction objective, trained on a corpus of human communication, has implicitly encoded emotional cues as high-dimensional features that strongly condition model output. Thus, emotional labels serve as a deterministic steering mechanism within the model’s probabilistic generation space where emotional cues (e.g., “frustration,” “satisfaction”) are strongly correlated with specific stances and potential actions. This makes an emotional signal a available control mechanism, which can directly shape an opponent LLM’s response generation.
@@
Despite this, current LLM agents are fundamentally limited in their use of emotion. Techniques like Direct Preference Optimization (DPO) (gao2025emo) and Reinforcement Learning from Human Feedback (RLHF) (kasbouya2025emotional) have made them proficient in recognizing and reacting to human emotional cues, but these capabilities remain reactive. They fail to leverage emotion proactively as a tool to influence the negotiation trajectory and assert strategic dominance. This leaves them vulnerable in adversarial LLM-vs-LLM interactions, where they exhibit three critical deficiencies:
@@ 2.1. Emotions in Negotiation Dynamics
While personality traits influence general negotiation tendencies (e.g., agreeableness correlating with cooperative strategies) (huang2024personality), emotions dominate real-time bargaining dynamics through three key mechanisms. First, temporal alignment ensures that negotiations—unfolding over rapid conversational turns lasting seconds to minutes—operate on a timescale closely aligned with emotional fluctuations, such as facial expressions and transient mood shifts, thereby often overriding longer-term personality influences (griessmair2015emotions). Second, strategic flexibility allows negotiators to adapt emotional expressions contextually; for instance, displaying tactical anger to convey firmness or simulate scarcity, temporarily overriding personality defaults (cite). Third, interactive amplification describes the process of emotional contagion, where one party’s affective stance elicits corresponding emotions in the opponent, forming feedback loops that modulate bargaining power over multiple turns (olekalns2014feeling). This fine-grained affective interplay underpins dynamic shifts in human negotiation behavior, making emotions a decisive factor in real-time decision-making.
@@ 2.2. LLM Agents in Negotiation Systems
Current LLM-based negotiation systems demonstrate significant capabilities in emotion recognition but fall short in strategic emotional adaptation. Frameworks like AgreeMate (chatterjee2024agreemate) and ACE (shea2024ace) show proficiency in inferring emotions through chain-of-thought reasoning, yet lack mechanisms for dynamic emotional strategy evolution. For instance, AgreeMate’s modular architecture optimizes fixed buyer/seller roles but cannot evolve emotional strategies amid negotiation. Similarly, ACE’s coaching system detects errors but cannot simulate the real-time emotional volatility that characterizes human bargaining. Consequently, current LLM agents remain largely reactive rather than proactive in their negotiation behavior.
Recent reinforcement learning advances have improved LLM negotiation capabilities. Methods such as Q-learning (watkins1992q) and PPO (schulman2017proximal) have been adapted to optimize negotiation policies via reward maximization. GENTEEL-NEGOTIATOR (priya2025genteel) further advances polite negotiation dialogue systems through mixture-of-expert reinforcement learning with specialized rewards for strategy, politeness, and coherence. Multi-agent evolutionary algorithms like EvoAgent (yuan2024evoagent) offer promising pathways for strategy enhancement through population-based optimization. However, these approaches face fundamental limitations in emotion-aware negotiation scenarios. Reinforcement learning methods struggle with the sparse-reward, high-volatility nature of negotiation environments, while evolutionary algorithms typically evaluate policies based on terminal outcomes, providing weak credit assignment for the nuanced interplay between consecutive emotional states in multi-turn negotiations. Crucially, these methods lack online learning capabilities essential for price negotiation across diverse products and scenarios. Real-world negotiation agents must adapt emotional strategies in real-time when facing different product types, price ranges, and debtor circumstances without requiring separate training for each variation. Our approach integrates evolutionary exploration with reinforcement learning to enable online emotion policy optimization, allowing agents to continuously adapt emotional strategies during execution across varied product negotiations.
@@ 4.1. Datasets
Likewise to huang2024personality, we selected a subset of negotiation cases from the CraigslistBargain dataset (he2018decoupling)—a widely-used benchmark for negotiation studies—for evaluation. This subset comprises 20 distinct multi-turn negotiation scenarios spanning diverse product categories, including electronics, furniture, vehicles, and housing. Each scenario specifies: (1) product details (name, category, and description), (2) agent-specific target prices for the seller, and (3) emotion annotations that support realistic bargaining dynamics. The selected dataset spans a wide price range (\$50–\$5,000) and includes varied item conditions (brand new or used), thereby enabling comprehensive evaluation of negotiation policies under heterogeneous market conditions.